# The Identification Power of

### v1

@@ Appendix C Proofs
Taking expectations of (43) and (44) lead to
@@
Hence, it follows that
$\theta^{U}-\theta_{U}$ $=\frac{1}{2}\int\int(\mathbb{E}[A_{W}^{\ast}(X)]-\mathbb{E}[A_{W}^{\ast}(S,X)])d\psi_{c}(y_{1},y_{0})$
$=\frac{1}{2}\int\int(F_{(+)}^{\ast}(y_{1}\vee y_{0},y_{1}\wedge y_{0})-F^{\ast,(+)}(y_{1}\vee y_{0},y_{1}\wedge y_{0}))d\psi_{c}(y_{1},y_{0})$
$+\frac{1}{2}\int\int(F_{(+)}^{\ast}(y_{1}\wedge y_{0},y_{1}\vee y_{0})-F^{\ast,(+)}(y_{1}\wedge y_{0},y_{1}\vee y_{0}))d\psi_{c}(y_{1},y_{0}),$ (45)
where $\theta^{U}-\theta_{U}\geq 0$ holds from the inequality (42). Therefore, if $\psi(\cdot,\cdot)$ is strict super-modular (implying that any rectangle in $(y_{1},y_{0})$-plane has a positive $\psi_{c}$ measure), it follows from (42) and (45) that $\theta^{U}=\theta_{U}$ iff $F^{\ast,(+)}(y_{1},y_{0})=F_{(+)}^{\ast}(y_{1},y_{0})$ for $\psi_{c}$-almost all $(y_{1},y_{0})$. Furthermore, for $\psi_{c}$-almost every $(y_{1},y_{0})$, $F^{\ast,(+)}(y_{1},y_{0})=F_{(+)}^{\ast}(y_{1},y_{0})$ holds iff
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1>0\right)\in\{0,1\}.$
@@
$2\theta_{L}$ $=\mathbb{E}[\psi(Y_{1},Y_{1})]+\mathbb{E}[\psi(Y_{0},Y_{0})]-\int\int\mathbb{E}[A_{M}^{\ast}(S,X)]d\psi_{c}(y_{1},y_{0}).$
@@
$\theta_{L}-\theta^{L}$ $=\frac{1}{2}\int\int(\mathbb{E}[A_{M}^{\ast}(S,X)]-\mathbb{E}[A_{M}^{\ast}(X)])d\psi_{c}(y_{1},y_{0})$
$=\frac{1}{2}\int\int(F^{\ast,(-)}(y_{1}\vee y_{0},y_{1}\wedge y_{0})-F_{(-)}^{\ast}(y_{1}\vee y_{0},y_{1}\wedge y_{0}))d\psi_{c}(y_{1},y_{0})$
$+\frac{1}{2}\int\int(F^{\ast,(-)}(y_{1}\wedge y_{0},y_{1}\vee y_{0})-F_{(-)}^{\ast}(y_{1}\wedge y_{0},y_{1}\vee y_{0}))d\psi_{c}(y_{1},y_{0}),$ (46)
where $\theta_{L}-\theta^{L}\geq 0$ holds from the inequality (41). Hence, if $\psi(\cdot,\cdot)$ is strict super-modular (implying that any rectangle in $(y_{1},y_{0})$-plane has a positive $\psi_{c}$ measure), it follows from (41) and (46) that $\theta^{L}=\theta_{L}$ iff $F^{\ast,(-)}(y_{1},y_{0})=F_{(-)}^{\ast}(y_{1},y_{0})$ for $\psi_{c}$-almost all $(y_{1},y_{0})$. Furthermore, for $\psi_{c}$-almost every $(y_{1},y_{0})$, $F^{\ast,(-)}(y_{1},y_{0})=F_{(-)}^{\ast}(y_{1},y_{0})$ holds iff
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)-F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)<0\right)\in\{0,1\}.$
We have so far shown the result under condition (a).
@@
Taking exceptions of (47) and (48) leads to
@@
where $\theta_{L}-\theta^{L}\geq 0$ holds from (41). Hence, if $\psi(\cdot,\cdot)$ is strict super-modular (implying that any rectangle in $(y_{1},y_{0})$-plane has a positive $\psi_{c}$ measure), it follows from (41) and (49) that $\theta^{L}=\theta_{L}$ iff $F^{\ast,(-)}(y_{1},y_{0})=F_{(-)}^{\ast}(y_{1},y_{0})$ for $\psi_{c}$-almost all $(y_{1},y_{0})$. Furthermore, for $\psi_{c}$-almost every $(y_{1},y_{0})$, $F^{\ast,(-)}(y_{1},y_{0})=F_{(-)}^{\ast}(y_{1},y_{0})$ holds iff $\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)-F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)<0\right)\in\{0,1\}.$
For the upper bounds under condition (b), by the similar argument, we can show that $\theta^{U}=\theta_{U}$ iff $\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1>0\right)\in\{0,1\}.$ We have then shown the result under condition (a). ∎
@@
Fix $\delta$. We first show part (i). Under the stated conditions and the definition of $\widetilde{\Theta}_{I}$, Theorem 3.5 (i) in Fan et al. (2017) implies that $\widetilde{\Theta}_{I}=[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)]$. Proposition 3.1 then establishes that $\Theta_{I}=\widetilde{\Theta}_{I}=[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)]$.
We now turn to part (ii). Under the stated conditions and the definition of $\widetilde{\Theta}_{IC}$, applying Theorem 3.5 (i) in Fan et al. (2017) with $X$ replaced by $(S,X)$ yields $\widetilde{\Theta}_{IC}=[F_{L,\varphi}(\delta),F_{U,\varphi}(\delta)]$. Theorem 3.2 then implies that $\Theta_{IC}=\widetilde{\Theta}_{IC}=[F_{L,\varphi}(\delta),F_{U,\varphi}(\delta)]$. ∎
@@
We provide a proof for the lower bounds. The proof for the upper bounds is similar and thus omitted. By definitions of $F_{\varphi}^{L}(\delta)$ and $F_{L,\varphi}(\delta)$ and Jensen’s inequality, we obtain:
@@
$=F_{L,\varphi}(\delta).$
Note that $Y_{d}$ ($d=0,1$) are assumed to be continuous random variables. Then, $F_{Y_{d}|SX}^{\ast}(y|S,X)$ and $F_{Y_{d}|X}^{\ast}(y|X)$ are continuous and thus $\sup_{y\in\mathcal{Y}_{1}}F_{Y_{1}|SX}^{\ast}(y|S,X)=1$ and $\sup_{y\in\mathcal{Y}_{1}}F_{Y_{1}|X}^{\ast}(y|X)=1$, implying $\sup_{y\in\mathcal{Y}_{1}}\{F_{Y_{1}|SX}^{\ast}(y|S,X)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|S,X)-1\}\geq 0$ and $\sup_{y\in\mathcal{Y}_{1}}\{F_{Y_{1}|X}^{\ast}(y|X)+F_{Y_{0}|X}^{\ast}(\tilde{\varphi}_{y}(\delta)|X)-1\}\geq 0$ a.s. Therefore, the inequality above becomes
@@
Let $G_{\varphi}(y,s,x)=F_{Y_{1}|SX}^{\ast}(y|s,x)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|s,x)-1$. Then, $\sup_{y\in\mathcal{Y}_{1}}\mathbb{E}[G(y,S,X)]=\mathbb{E}[G(\bar{y},S,X)]$ and it follows from (50) that $F_{\varphi}^{L}(\delta)=F_{L,\varphi}(\delta)$ iff $\mathbb{E}[\sup_{y\in\mathcal{Y}_{1}}G(y,S,X)]=\mathbb{E}[G(\bar{y},S,X)]$. Since $\sup_{y\in\mathcal{Y}_{1}}G(y,s,x)\geq G(\bar{y},s,x)$ for all $(s,x)\in\{0,1\}\times\mathcal{X}$, it implies that $\mathbb{E}[\sup_{y\in\mathcal{Y}_{1}}G(y,S,X)]=\mathbb{E}[G(\bar{y},S,X)]$ iff $\sup_{y\in\mathcal{Y}_{1}}G(y,s,x)=G(\bar{y},s,x)$ for almost all $(s,x)\in\{0,1\}\times\mathcal{X}$; that is, $G_{\varphi}(y,s,x)$ reaches its maximum value uniformly at $\bar{y}$ for almost all $(s,x)\in\{0,1\}\times\mathcal{X}$. ∎
@@
Let
@@
and define $\widetilde{\Theta}_{IC}^{\dagger}\equiv\left\{\mathbb{E}_{F_{Y_{1}Y_{0}SX}}[\psi(Y_{1},Y_{0})]:F_{Y_{1}Y_{0}SX}\in\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}\right\}$. We also define a functional $\Gamma:\mathcal{F}_{Y_{1}Y_{0}SX}\rightarrow\mathbb{R}$ as $\Gamma(F_{Y_{1}Y_{0}SX})\equiv\mathbb{E}_{F_{Y_{1}Y_{0}SX}}[\psi(Y_{1},Y_{0})]$. Note that $\widetilde{\Theta}_{IC}^{\dagger}=\left\{\Gamma(F_{Y_{1}Y_{0}SX}):F_{Y_{1}Y_{0}SX}\in\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}\right\}$.
Since all restrictions (14)–(16) are linear and $\mathcal{F}_{Y_{1}Y_{0}SX}$ is a convex set, $\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}$ is convex as well. Then since $\Gamma:\mathcal{F}_{Y_{1}Y_{0}SX}\rightarrow\mathbb{R}$ is linear, it follows from convexity of $\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}$ that $\widetilde{\Theta}_{IC}^{\dagger}$ is a convex set, and so its closure is
@@
We next show that $\widetilde{\Theta}_{IC}^{\dagger}=\Theta_{IC}^{\dagger}$. Regarding conditional joint distributions $F_{Y_{1}Y_{0}|SX}$, we define the following conditions for all $(s,x)$:
@@
$\mathbb{P}_{F_{Y_{1}Y_{0}|SX}}(Y_{1}-Y_{0}>c\mid S=1,X=x)\geq\mathbb{P}_{F_{Y_{1}Y_{0}|SX}}(Y_{1}-Y_{0}>c\mid S=0,X=x)\mbox{ for all }c$ (52)
For all $x\in\mathcal{X}$, we define the following class of pairs of conditional copula functions:
@@
This is a class of pairs of conditional copula functions that replicate the conditional joint distributions $F_{Y_{1}Y_{0}|SX}$ satisfying conditions (51) and (52). Then $\widetilde{\Theta}_{IC}^{\dagger}$ can be expressed as
$\widetilde{\Theta}_{IC}^{\dagger}$ $=\left\{\theta:\theta=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)|S,X)\right]\right$.
@@
We begin by showing that $\Theta_{IC}^{\dagger}\subseteq\widetilde{\Theta}_{IC}^{\dagger}$. Fix any $\theta\in\Theta_{IC}^{\dagger}$. By definition of $\Theta_{IC}^{\dagger}$, there exists a distribution function $F\in\widetilde{\mathcal{F}}^{\ast}$ such that $\theta=\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$. Since $F_{YDGX}=F_{YDGX}^{\ast}$ holds and $F$ satisfies Assumptions 2.12.5, Lemma 3.1 implies that $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$ for $d=0,1$.
Because $F_{Y_{1}Y_{0}SX}\in\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}$, which implies $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$, by Sklar’s theorem and definition of $\widetilde{\mathcal{C}}_{x}$, there exists a pair of conditional copula functions $(C(\cdot,\cdot|0,x),C(\cdot,\cdot|1,x))\in\widetilde{\mathcal{C}}_{x}$ that reproduces $F_{Y_{1}Y_{0}|SX}$ as
@@
Then it follows that
@@
where the fourth line follows from $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$, and the last line follows from (53). We thus have $\theta\in\widetilde{\Theta}_{IC}^{\dagger}$. Since this discussion holds for any $\Theta_{IC}^{\dagger}$, we have $\Theta_{IC}^{\dagger}\subseteq\widetilde{\Theta}_{IC}^{\dagger}$.
We next show that $\widetilde{\Theta}_{IC}^{\dagger}\subseteq\Theta_{IC}^{\dagger}$. Fix any $\theta\in\widetilde{\Theta}_{IC}^{\dagger}$. By definition of $\widetilde{\Theta}_{IC}^{\dagger}$, there exists a pair of conditional copula functions $(C(\cdot,\cdot|s,x),C(\cdot,\cdot|s,x))\in\widetilde{\mathcal{C}}$ such that
We will show that there exists a distribution function $F$ in $\widetilde{\mathcal{F}}^{\ast}$ that replicates $\theta$ as $\theta=\mathbb{E}_{F}\left[\psi(Y_{1},Y_{0})\right]$. Specifically, we construct such a a distribution function $F$ of $(Y_{1},Y_{0},Y,D,S,G,X)$ in the following way:
$F=F_{Y|Y_{1}Y_{0}DSGX}\times F_{S|Y_{1}Y_{0}DGX}\times F_{Y_{1}Y_{0}|DGX}\times F_{DGX},$
where

### v4

@@ Appendix C Proofs and a Preliminary Lemma
Taking expectations of (43) and (44) yields
@@
Hence,
$\theta^{U}-\theta_{U}$ $=\frac{1}{2}\int\int(\mathbb{E}[A_{W}^{\ast}(S,X)]-\mathbb{E}[A_{W}^{\ast}(X)])d\psi_{c}(y_{1},y_{0})$
$=\frac{1}{2}\int\int(F^{\ast,(+)}(y_{1}\vee y_{0},y_{1}\wedge y_{0})-F_{(+)}^{\ast}(y_{1}\vee y_{0},y_{1}\wedge y_{0}))d\psi_{c}(y_{1},y_{0})$
$+\frac{1}{2}\int\int(F^{\ast,(+)}(y_{1}\wedge y_{0},y_{1}\vee y_{0})-F_{(+)}^{\ast}(y_{1}\wedge y_{0},y_{1}\vee y_{0}))d\psi_{c}(y_{1},y_{0}),$ (45)
where $\theta^{U}-\theta_{U}\geq 0$ holds by (42).
Hence, if $\psi(\cdot,\cdot)$ is strict super-modular (implying that any rectangle in $(y_{1},y_{0})$-plane has a positive $\psi_{c}$ measure), it follows from (42) and (45) that $\theta^{U}=\theta_{U}$ if and only if $F^{\ast,(+)}(y_{1},y_{0})=F_{(+)}^{\ast}(y_{1},y_{0})$ for $\psi_{c}$-almost all $(y_{1},y_{0})$. Moreover, from equation (42), it follows that for $\psi_{c}$-almost every $(y_{1},y_{0})$, the condition $F^{\ast,(+)}(y_{1},y_{0})=F_{(+)}^{\ast}(y_{1},y_{0})$ holds if and only if
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1>0\middle|X\right)\in\{0,1\}\text{ a.s.}$
@@
$2\theta_{L}$ $=\mathbb{E}[\psi(Y_{1},Y_{1})]+\mathbb{E}[\psi(Y_{0},Y_{0})]-\int\int\mathbb{E}[A_{M}^{\ast}(S,X)]d\psi_{c}(y_{1},y_{0}),$
@@
$\theta_{L}-\theta^{L}$ $=\frac{1}{2}\int\int(\mathbb{E}[A_{M}^{\ast}(X)]-\mathbb{E}[A_{M}^{\ast}(S,X)])d\psi_{c}(y_{1},y_{0})$
$=\frac{1}{2}\int\int(F_{(-)}^{\ast}(y_{1}\vee y_{0},y_{1}\wedge y_{0})-F^{\ast,(-)}(y_{1}\vee y_{0},y_{1}\wedge y_{0}))d\psi_{c}(y_{1},y_{0})$
$+\frac{1}{2}\int\int(F_{(-)}^{\ast}(y_{1}\wedge y_{0},y_{1}\vee y_{0})-F^{\ast,(-)}(y_{1}\wedge y_{0},y_{1}\vee y_{0}))d\psi_{c}(y_{1},y_{0}),$ (46)
where $\theta_{L}-\theta^{L}\geq 0$ holds by (41).
Hence, if $\psi(\cdot,\cdot)$ is strict super-modular (implying that any rectangle in $(y_{1},y_{0})$-plane has a positive $\psi_{c}$ measure), it follows from (41) and (46) that $\theta^{L}=\theta_{L}$ if and only if $F^{\ast,(-)}(y_{1},y_{0})=F_{(-)}^{\ast}(y_{1},y_{0})$ for $\psi_{c}$-almost all $(y_{1},y_{0})$. Furthermore, from equation (41), it follows that for $\psi_{c}$-almost every $(y_{1},y_{0})$, the condition $F^{\ast,(-)}(y_{1},y_{0})=F_{(-)}^{\ast}(y_{1},y_{0})$ holds if and only if
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)-F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)<0\middle|X\right)\in\{0,1\}\text{ a.s.}$
We have thus established the result under condition (a).
@@
Taking exceptions of (47) and (48) yields
@@
where $\theta_{L}-\theta^{L}\geq 0$ holds from (41).
Hence, if $\psi(\cdot,\cdot)$ is strict super-modular (implying that any rectangle in $(y_{1},y_{0})$-plane has a positive $\psi_{c}$ measure), it follows from (41) and (49) that $\theta^{L}=\theta_{L}$ if and only if $F^{\ast,(-)}(y_{1},y_{0})=F_{(-)}^{\ast}(y_{1},y_{0})$ for $\psi_{c}$-almost all $(y_{1},y_{0})$. Furthermore, for $\psi_{c}$-almost every $(y_{1},y_{0})$, the condition $F^{\ast,(-)}(y_{1},y_{0})=F_{(-)}^{\ast}(y_{1},y_{0})$ holds if and only if
$\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)-F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)<0\middle|X\right)\in\{0,1\}\text{ a.s.}$
For the upper bounds under condition (b), by a similar argument, we can show that $\theta^{U}=\theta_{U}$ if and only if $\mathbb{P}\left(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X)+F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)-1>0\middle|X\right)\in\{0,1\}$ a.s. Thus, the result under condition (b) has now been established. ∎
@@
Fix $\delta$. We first prove part (i). Under the stated conditions and by the definition of $\widetilde{\Theta}_{I}$, Theorem 3.5(i) in Fan et al. (2017) implies $\widetilde{\Theta}_{I}=[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)]$. Proposition 3.1 then yields $\Theta_{I}=\widetilde{\Theta}_{I}=[F_{\varphi}^{L}(\delta),F_{\varphi}^{U}(\delta)]$.
For part (ii), under the stated conditions and by the definition of $\widetilde{\Theta}_{IC}$, applying Theorem 3.5(i) in Fan et al. (2017) with $X$ replaced by $(S,X)$ gives $\widetilde{\Theta}_{IC}=[F_{L,\varphi}(\delta),F_{U,\varphi}(\delta)]$. Theorem 3.3 then yields $\Theta_{IC}=\widetilde{\Theta}_{IC}=[F_{L,\varphi}(\delta),F_{U,\varphi}(\delta)]$. ∎
@@
We provide a proof for the lower bounds. The proof for the upper bounds is analogous and therefore omitted. By the definitions of $F_{\varphi}^{L}(\delta)$ and $F_{L,\varphi}(\delta)$ and by Jensen’s inequality, we have
@@
$=F_{L,\varphi}(\delta).$ (50)
Note that $Y_{d}$ ($d=0,1$) are assumed to be continuous random variables. Hence $F_{Y_{d}|SX}^{\ast}(y|S,X)$ and $F_{Y_{d}|X}^{\ast}(y|X)$ are continuous, and thus $\sup_{y\in\mathcal{Y}_{1}}F_{Y_{1}|SX}^{\ast}(y|S,X)=1$ and $\sup_{y\in\mathcal{Y}_{1}}F_{Y_{1}|X}^{\ast}(y|X)=1$. This implies that
$\sup_{y\in\mathcal{Y}_{1}}\{F_{Y_{1}|SX}^{\ast}(y|S,X)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|S,X)-1\}\geq 0\mbox{ a.s.},$
and
$\sup_{y\in\mathcal{Y}_{1}}\{F_{Y_{1}|X}^{\ast}(y|X)+F_{Y_{0}|X}^{\ast}(\tilde{\varphi}_{y}(\delta)|X)-1\}\geq 0\mbox{ a.s.}$
Therefore, equation (50) simplifies to
@@
Let $G_{\varphi}(y,s,x)=F_{Y_{1}|SX}^{\ast}(y|s,x)+F_{Y_{0}|SX}^{\ast}(\tilde{\varphi}_{y}(\delta)|s,x)-1$. Then $\sup_{y\in\mathcal{Y}_{1}}\mathbb{E}[G(y,S,X)|X=x]=\mathbb{E}[G(\bar{y}(x),S,X)|X=x]$ and it follows from (51) that $F_{\varphi}^{L}(\delta)=F_{L,\varphi}(\delta)$ if and only if $\mathbb{E}[\sup_{y\in\mathcal{Y}_{1}}G(y,S,X)|X=x]=\mathbb{E}[G(\bar{y}(x),S,X)|X=x]$ for almost all $x\in\mathcal{X}$.
Since $\sup_{y\in\mathcal{Y}_{1}}G(y,s,x)\geq G(\bar{y}(x),s,x)$ for all $(s,x)\in\{0,1\}\times\mathcal{X}$, it follows that
$\mathbb{E}[\sup_{y\in\mathcal{Y}_{1}}G(y,S,X)|X=x]=\mathbb{E}[G(\bar{y}(x),S,X)|X=x]$
for almost all $x\in\mathcal{X}$ if and only if $\sup_{y\in\mathcal{Y}_{1}}G(y,s,x)=G(\bar{y}(x),s,x)$ for almost all $(s,x)\in\{0,1\}\times\mathcal{X}$; that is, $G_{\varphi}(y,0,x)$ and $G_{\varphi}(y,1,x)$ attain their maxima at the common point $\bar{y}(x)$ for almost all $x\in\mathcal{X}$. ∎
@@
Let $\mathcal{F}$ denote the class of all distribution functions of the variables $(Y_{1},Y_{0},Y,D,S,G,X)$. For any generic subclass $\widetilde{\mathcal{F}}\subseteq\mathcal{F}$, let $\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}=\{F_{Y_{1}Y_{0}SX}:F\in\widetilde{\mathcal{F}}\}$ be the corresponding set of marginal distributions of $(Y_{1},Y_{0},S,X)$. Next, let $\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}$ denote the set of distribution functions of $(Y_{1},Y_{0},S,X)$ that satisfy constraints (16)–(18):
@@
Define
$\widetilde{\Theta}_{IC}^{\dagger}\equiv\bigl\{\mathbb{E}_{F_{Y_{1}Y_{0}SX}}[\psi(Y_{1},Y_{0})]:F_{Y_{1}Y_{0}SX}\in\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}\bigr\}.$
We also introduce the functional $\Gamma:\mathcal{F}_{Y_{1}Y_{0}SX}\rightarrow\mathbb{R}$ defined by $\Gamma(F_{Y_{1}Y_{0}SX})\equiv\mathbb{E}_{F_{Y_{1}Y_{0}SX}}[\psi(Y_{1},Y_{0})]$. It then follows that $\widetilde{\Theta}_{IC}^{\dagger}=\bigl\{\Gamma(F_{Y_{1}Y_{0}SX}):F_{Y_{1}Y_{0}SX}\in\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}\bigr\}.$
Since all restrictions (16)–(18) are linear and $\mathcal{F}_{Y_{1}Y_{0}SX}$ is convex, $\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}$ is convex as well. Moreover, because $\Gamma:\mathcal{F}_{Y_{1}Y_{0}SX}\rightarrow\mathbb{R}$ is linear, the convexity of $\widetilde{\mathcal{F}}_{Y_{1}Y_{0}SX}^{\dagger}$ implies that $\widetilde{\Theta}_{IC}^{\dagger}$ is convex. Hence its closure is
@@
We next show that $\widetilde{\Theta}_{IC}^{\dagger}=\Theta_{IC}^{\dagger}$. For conditional joint distributions $F_{Y_{1}Y_{0}|SX}$, we consider the following conditions for all $(s,x)$:
$F_{Y_{d}|Y_{d^{\prime}}\leq y,S=s,X=x}(t)\geq F_{Y_{d}|Y_{d^{\prime}}\leq y^{\prime},S=s,X=x}(t)$
@@
$\mathbb{P}_{F_{Y_{1}Y_{0}|SX}}(Y_{1}-Y_{0}>c\mid S=1,X=x)\geq\mathbb{P}_{F_{Y_{1}Y_{0}|SX}}(Y_{1}-Y_{0}>c\mid S=0,X=x)\mbox{ for all }c\in\mathbb{R}$ (53)
Conditions (52) and (53) are equivalent to Assumptions 5.1 and 5.2, respectively, conditional on $(S,X)=(s,x)$.
For each $x\in\mathcal{X}$, we define the following class of pairs of conditional copula functions:
@@
That is, $\widetilde{\mathcal{C}}_{x}$ consists of all pairs of conditional copula functions that generate conditional joint distributions $F_{Y_{1}Y_{0}|SX}$ satisfying conditions (52) and (53). Note that $\widetilde{\mathcal{C}}_{x}$ is nonempty, since there exists a pair $(C^{\ast}(\cdot,\cdot|0,x),C^{\ast}(\cdot,\cdot|1,x))\in\widetilde{\mathcal{C}}_{x}$ such that
$F_{Y_{1}Y_{0}|SX}^{\ast}(\cdot,\cdot|s,x)=C^{\ast}(F_{Y_{1}|SX}^{\ast}(\cdot|s,x),F_{Y_{0}|SX}^{\ast}(\cdot|s,x)),$
for all $(s,x)$.
Then $\widetilde{\Theta}_{IC}^{\dagger}$ can be expressed as
$\widetilde{\Theta}_{IC}^{\dagger}$ $=\left\{\theta:\theta=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)|S,X\bigr)\right]\right$.
@@
We first show that $\Theta_{IC}^{\dagger}\subseteq\widetilde{\Theta}_{IC}^{\dagger}$. Fix any $\theta\in\Theta_{IC}^{\dagger}$. By the definition of $\Theta_{IC}^{\dagger}$, there exists a distribution function $F\in\widetilde{\mathcal{F}}^{\ast}$ such that $\theta=\mathbb{E}_{F}[\psi(Y_{1},Y_{0})]$. Since $F_{YDGX}=F_{YDGX}^{\ast}$ and $F$ satisfies Assumptions 2.12.5, Lemma 3.2 implies that $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$ for $d=0,1$.
By Sklar’s theorem and the definition of $\widetilde{\mathcal{C}}_{x}$, the equalities $F_{Y_{d}\mid SX}=F_{Y_{d}\mid SX}^{\ast}$ for $d=0,1$ guarantee the existence of a pair of conditional copula functions $(C(\cdot,\cdot\mid 0,x),C(\cdot,\cdot\mid 1,x))\in\widetilde{\mathcal{C}}_{x}$ such that
@@
It then follows that
@@
where the fourth line uses $F_{Y_{d}SX}=F_{Y_{d}SX}^{\ast}$, and the last line follows from (54). Thus, we have $\theta\in\widetilde{\Theta}_{IC}^{\dagger}$. Since this argument holds for any $\theta\in\Theta_{IC}^{\dagger}$, it follows that $\Theta_{IC}^{\dagger}\subseteq\widetilde{\Theta}_{IC}^{\dagger}$.
We next show that $\widetilde{\Theta}_{IC}^{\dagger}\subseteq\Theta_{IC}^{\dagger}$. Fix any $\theta\in\widetilde{\Theta}_{IC}^{\dagger}$. By the definition of $\widetilde{\Theta}_{IC}^{\dagger}$, there exists a pair of conditional copula functions $(C(\cdot,\cdot|s,x),C(\cdot,\cdot|s,x))\in\widetilde{\mathcal{C}}_{x}$ such that
$\theta=\mathbb{E}_{F_{SX}^{\ast}}\left[\int\int\psi(y_{1},y_{0})dC\bigl(F_{Y_{1}|SX}^{\ast}(y_{1}|S,X),F_{Y_{0}|SX}^{\ast}(y_{0}|S,X)\big|S,X\bigr)\right].$
We now show that there exists a distribution function $F\in\widetilde{\mathcal{F}}^{\ast}$ that reproduces $\theta$ as $\theta=\mathbb{E}_{F}\left[\psi(Y_{1},Y_{0})\right]$. Specifically, we construct such a distribution function $F$ of $(Y_{1},Y_{0},Y,D,S,G,X)$ hierarchically, defined by